OpenStudy (anonymous):

solve the equation.check your answer.m/20=5/4

thomaster (thomaster):

\(\sf\large\dfrac{m}{20}=\dfrac{5}{4}\) multiply both sides by 20 and you'll get this: \(\sf\large m=\dfrac{5*20}{4}\)